# Examining dark neutron decay using f-mode gravitational waves

### v2

@@ 9 Conclusion
In this work, we investigated the impact of nucleonic, strange matter, and dark matter equations of state on the f-mode oscillation properties of neutron stars, with particular attention to damping times and universal relations. Our results demonstrate that the nucleon-only EoS is consistently the stiffest, while the inclusion of dark matter generally softens the stellar structure, unless significant DM–DM self-interactions are present. Increasing the self-repulsion parameter $G$ suppresses dark matter production, thereby stiffening the EoS and bringing the neutron-decay scenario closer to the nucleonic case. In contrast, strange matter exhibits longer damping times and notable deviations at low densities, reflecting its softer pressure support relative to nucleonic matter.
Importantly, we find that neutron decay into dark matter leads to stellar configurations that remain compatible with current astrophysical constraints. Moreover, the corresponding f-mode frequencies of such stars fall squarely within the detection range of modern gravitational-wave interferometers, underscoring their observational relevance.
Despite these compositional differences, the mass-scaled angular frequency retains quasi-universal behavior when expressed as a function of compactness ($C=M/R$) and the structural parameter $\psi=\sqrt{M^{3}/I}$. This universality is preserved even in the presence of exotic matter, with only minor deviations in the case of strange matter and non-self-interacting dark matter. The persistence of such relations highlights their robustness and utility as probes of neutron star structure.
Astrophysically, these results suggest that f-mode oscillations and their associated gravitational-wave signals can provide meaningful constraints on neutron star properties without requiring precise knowledge of the microscopic EoS. At the same time, the systematic deviations observed in exotic matter scenarios offer potential observational signatures that could help discriminate between standard nucleonic models and stars containing dark or strange matter. Future detections of f-mode gravitational waves by next-generation interferometers could therefore serve as a powerful testbed for probing the presence of exotic matter in compact stars and for constraining the strength of dark matter self-interactions. In particular, estimating the signal-to-noise ratios of the predicted f-modes in next-generation detectors such as the Einstein Telescope and Cosmic Explorer, for realistic excitation energies and astrophysical channels (post-merger remnants, proto-neutron stars, or glitches), will be an important direction. Such studies could establish whether the parameter space explored here can be probed observationally, and whether stacking strategies or Bayesian inference using universal relations can enhance detectability. This will be pursued in future work.
Acknowledgments

### v3

@@ 9 Conclusion
In this work, the f-mode oscillation properties have been explored of neutron stars containing nucleonic matter, strange matter, and dark matter produced through neutron decay, with particular attention to the role of dark matter self-interactions. By computing f-mode frequencies, damping times, characteristic strains, and signal-to-noise ratios, the detectability of these oscillations with the Einstein Telescope (ET) is assessed.
The nucleonic EoS remains the stiffest, producing systematically higher f-mode frequencies, while the inclusion of dark matter generally softens the stellar structure. Strong DM–DM self-repulsion ($G\gtrsim 26$ fm2) counteracts this softening and restores compatibility with astrophysical mass constraints. Strange matter predicts longer damping times and lower frequencies, shifting the signal toward the ET’s most sensitive frequency band.
For an assumed gravitational-wave energy release of $E=10^{52}$ erg, damping time = 0.1 sec and a source distance of $D=25$ Mpc, the characteristic strain and SNR of self-interacting dark matter and strange matter models exceed the ET sensitivity threshold below $\sim 2.1$ kHz. In contrast, the non–self-interacting DM case (G0) lies below the sensitivity curve, suggesting that such signals would be undetectable.
These results indicate that third-generation gravitational-wave observatories could realistically probe exotic matter scenarios through f-mode asteroseismology. A confirmed detection of f-mode oscillations in the $\sim$1–2 kHz range would provide strong evidence for exotic matter in neutron star interiors, whereas a lack of such signals could place meaningful constraints on dark matter self-interaction strengths and the viability of strange matter.
This study demonstrates that f-mode oscillations represent a promising observational probe of neutron star interiors and exotic matter physics. The persistence of universal relations across different EoS models highlights their robustness, while the deviations introduced by strange matter and non-self-interacting dark matter suggest possible observational discriminants. Future detections of f-mode gravitational waves by instruments such as ET and Cosmic Explorer could therefore open a new window into the microphysics of dense matter, providing unprecedented insight into the role of dark matter and exotic phases in compact stars. This analysis extends previous work by explicitly connecting exotic matter EoS models to gravitational-wave observability, making this the first quantitative detectability study of neutron-decay dark matter through f-mode oscillations. The results indicate that third-generation interferometers such as ET could provide decisive tests of dark matter self-interactions and the presence of exotic matter in compact stars.
Acknowledgment